# Unusual switch from low-temperature T-quadratic resistivity in the underdoped pseudogap phase of cuprate superconductors to low-temperature T-linear resistivity in the overdoped strange-metal phase

### v1

@@ II.2 Pseudogap in charge and spin channels
Based on the full spin propagator (17) [then the dynamical spin structure factor], the dynamical spin response of cuprate superconductors in the normal-state has been investigatedKuang15 ; Yuan01 ; Feng98 , where the obtained results show the existence of damped but well-defined dispersive spin excitations in the whole doping phase diagram. Moreover, the low-energy spin fluctuation is dominated by the process from the mobile charge carriers, while the high-energy spin excitation on the other hand retains roughly constant energy as a function of doping, with spectral weight and dispersion relation comparable to those in the corresponding SC-state. All these results are qualitatively consistent with the experimental observations Fujita12 ; Birgeneau89 ; Fong95 ; Yamada98 ; Arai99 ; Bourges00 ; He01 ; Hayden04 ; Tranquada04 ; Bourges05 .
@@ II.3 Electronic structure in the underdoped pseudogap state
For the understanding of the nature of the electronic structure in cuprate superconductors, it is needed to derive the electron propagator, which is characterized by the charge-spin recombination. In order to obtain the electron propagator, the full charge-spin recombination scheme has been developedFeng15a , where it has been shown that the coupling form between the electron and spin excitation in the $t$-$J$ model in the normal-state is the same as that between the charge carrier and spin excitation in Eq. (4), and then the self-consistent equation satisfied by the full electron propagator in the normal-state can be derived directly asFeng15a ,
where $G^{(0)}({\bf k},\omega)$ is the electron propagator of the $t$-$J$ model (1) in the tight-binding approximation, and has been derived as $G^{(0)-1}({\bf k},\omega)=\omega-\varepsilon_{\bf k}$ with the electron energy dispersion $\varepsilon_{\bf k}=-4t\gamma_{\bf k}+4t^{\prime}\gamma_{\bf k}^{\prime}+\mu$. In this case, the full electron propagator $G({\bf k},\omega)$ in Eq. (20) can be expressed explicitly as,
@@
with the electron normal self-energy,
$\times$ $P^{(0)}({\bf k},{\bf p},ip_{m}),~{}~{}~{}$
which has been evaluated asFeng15a ,
with the function,
where $\mu~{}(\nu)=1,~{}2$, $\bar{\varepsilon}_{\bf k}=Z_{\rm F}\varepsilon_{\bf k}$ is the renormalized electron energy dispersion, $Z^{-1}_{\rm F}=1-{\rm Re}\Sigma_{\rm pho}({\bf k},0)\mid_{{\bf k}=[\pi,0]}$ is the electron coherent weight, $\Sigma_{\rm pho}({\bf k},\omega)$ is the antisymmetric part of the electron normal self-energy $\Sigma_{\rm ph}({\bf k},\omega)$, $n^{(\nu)}_{{\rm 1B}{\bf pp^{\prime}}}=1+n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})+n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$, and $n^{(\nu)}_{{\rm 2B}{\bf pp^{\prime}}}=n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$. This electron normal self-energy characterizes a competition between the kinetic energy and magnetic energy in the $t$-$J$ model (1).
The electron normal self-energy (23) originated from the charge-carrier normal self-energy (10) is due to the charge-spin recombination, indicating that the normal-state pseudogap state originated from the charge-carrier pseudogap state is due to the charge-spin recombination. To see the nature of the normal-state pseudogap more clearly, the above electron normal self-energy in Eq. (23) can be rewritten as,
$\Sigma_{\rm ph}({\bf k},\omega)$ $=$ ${[\bar{\Delta}_{\rm PG}({\bf k})]^{2}\over\omega-\varepsilon_{0{\bf k}}},$ (25)
where $\varepsilon_{0{\bf k}}=L_{2{\bf k}}/L_{1{\bf k}}$ is the energy spectrum of $\Sigma_{\rm ph}({\bf k},\omega)$, and $\bar{\Delta}^{2}_{\rm PG}({\bf k})=L^{2}_{2{\bf k}}/L_{1{\bf k}}$ is the normal-state pseudogap, with the functions $L_{1}({\bf k})=-\Sigma_{\rm pho}({\bf k},\omega=0)$ and $L_{2}({\bf k})=-\Sigma_{\rm ph}({\bf k},\omega=0)$, while $\Sigma_{\rm pho}({\bf k},\omega=0)$ and the antisymmetric part $\Sigma_{\rm pho}({\bf k},\omega)$ of the electron normal self-energy can be derived directly from the electron normal self-energy $\Sigma_{\rm ph}({\bf k},\omega)$ in Eq. (23). This normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ is also identified as being a region of the electron normal self-energy in which the normal-state pseudogap anisotropically suppresses the electronic density of states on EFS. Since the normal-state pseudogap state is induced by the charge-carrier pseudogap state, the normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ [the normal-state pseudogap crossover temperature $T^{*}$] as a function of doping presents a similar behavior of the charge-carrier pseudogap $\bar{\Delta}^{({\rm h})}_{\rm pg}({\bf k})$ [the charge-carrier pseudogap crossover temperature $T^{*}_{\rm h}$]Feng15 ; Feng12 . This normal-state pseudogap crossover temperature $T^{*}$ is actually a crossover line below which a novel electronic state emerges, as exemplified by the presence of the Fermi arcs, the competing electronic orders, etc., and then the unconventional features of this novel electronic state can be well interpreted in terms of the formation of the normal-state pseudogap Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 .
@@
$I({\bf k},\omega)\propto n_{\rm F}(\omega)A({\bf k},\omega),~{}~{}~{}~{}$ (26)
@@
where ${\rm Re}\Sigma_{\rm ph}({\bf k},\omega)$ and ${\rm Im}\Sigma_{\rm ph}({\bf k},\omega)$ are the real and imaginary parts of the electron normal self-energy $\Sigma_{\rm ph}({\bf k},\omega)$, respectively. The electrons are renormalzed due to the electron scattering mediated by the spin excitation and then they acquire a finite lifetime.
The EFS topology is known to be crucial in the understanding of the unconventional superconductivity in cuprate superconductors Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 and the related anomalous normal-state properties Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 . The ARPES measured EFS can be described theoretically by the intensity map of the electron excitation spectrum (26) at zero energy $\omega=0$, where the locations of the EFS continuous contour in momentum space is determined directly by the poles of the electron spectral function (27),
@@
$\varepsilon_{\bf k}+{\rm Re}\Sigma_{\rm ph}({\bf k},0)=\bar{\varepsilon}_{\bf k}=0$, while the strong redistribution of the spectral weight on EFS is mainly governed by the momentum dependence of the normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ [then the single-electron scattering rate $\Gamma_{\bf k}(\omega)={\rm Im}\Sigma_{\rm ph}({\bf k},\omega)=\pi[\bar{\Delta}_{\rm PG}({\bf k})]^{2}\delta(\omega+\varepsilon_{0{\bf k}})]$. For a convenience in the following discussions of the electrical transport in the underdoped pseudogap phase, the EFS mapFeng16 ; Liu21 ; Cao21 ; Zeng22 at doping $\delta=0.09$ with temperature $T=0.002J$ is replotted in Fig. 3, where the BZ center has been shifted by [$\pi,\pi$], and AN, TFA, and ND indicate the antinode, tip of the Fermi arc, and node, respectively. It thus shows clearly that the antinodal region of EFS becomes partially gapped, leading to that EFS consists, not of a closed contour, but only of four disconnected Fermi arcs centered around the nodes, in qualitative agreement with the experimental results Loeser96 ; Norman98 ; Chatterjee06 ; Shi08 ; Sassa11 ; Kaminski15 ; Loret17 ; Loret18 ; He14 ; Comin14 ; Comin16 . However, the most of the spectral weight on the Fermi arcs locates at around the tips of the Fermi arcs, indicating that the electrons at around the tips of the Fermi arcs have a largest density of states Chatterjee06 ; Shi08 ; Sassa11 ; Kaminski15 ; Loret17 ; Loret18 ; He14 ; Comin14 ; Comin16 , and then charge order is driven by this EFS instability, with a characteristic wave vector corresponding to the tips of the Fermi arcsFeng16 ; Comin14 ; Comin16 .
The elementary excitations are parameterized by the electron spectral function (27), which has given a consistent description of the renormalization of the electrons in the underdoped pseudogap phaseFeng16 ; Liu21 ; Cao21 ; Zeng22 , where the obtained results show that the single-electron scattering rate has a well-pronounced peak structure at around the antinodal and nodal regions, which leads to the remarkable peak-dip-hump structure in the energy distribution curve. Moreover, the striking dispersion kink is induced by the inflection point in the single-electron scattering rate, while the spectral weight at around the dispersion kink is reduced highly by the corresponding peak in the real part of the electron normal self-energy. All these results are in qualitative agreement with the corresponding experimental results Dessau91 ; Randeria95 ; Fedorov99 ; Campuzano99 ; DMou17 ; Bogdanov00 ; Kaminski01 ; Johnson01 ; Iwasawa08 ; Plumb13 . More importantly, the anisotropic suppression of the electronic density of states on EFS by the normal-state pseudogap can affect the electrical transport in two ways Timusk99 : (i) through the reduction of the number of electron current-carrying states as a normal-state pseudogap forms; and (ii) since the electron current carriers are scattered by the spin excitation, through the reduction in the density of spin excitations.
@@ II.4 Electrical transport due to umklapp scattering from a spin excitation
We now turn to discuss the low-temperature electrical transport in the underdoped pseudogap phase. Theoretically, the Boltzmann transport equation is the cornerstone for the discussions of the electrical transportAbrikosov88 ; Mahan81 , since the Boltzmann transport equation is valid in the case of either the existence of the well-defined quasiparticles or the treatment of the electron interaction mediated by different bosonic modes within the Eliashberg approach. This builds on the following pioneering works: (i) in the early days of the electrical transport research in the conventional superconductorsPrange64 , a set of the transport equations has been derived in the Migdal’s approximation, where this coupled set of the transport equations for the electron and phonon distribution functions are correct even in the case of the absence of the well-defined quasiparticles. In particular, one of the forms of the electrical transport equation,
@@
In the following discussions, we study the the low-temperature electrical transport in the underdoped pseudogap phase based on the electrical Boltzmann transport equation (28). For the calculation of this electrical Boltzmann transport equation (28), the linear perturbation from the equilibrium in terms of the distribution function has been introduced Lee21 ; Ma23 ; Prange64 ,
@@
with the local shift of the chemical potential at a given patch of EFS $\tilde{\Phi}({\bf k})$, which satisfied the antisymmetric relation $\tilde{\Phi}(-{\bf k})=-\tilde{\Phi}({\bf k})$.
We substitute the above result in Eq. (29) into Eq. (28), and then linearize the electrical Boltzmann equation (28) as,
@@
The electron-electron collision term $I_{\rm e-e}$ in the electrical Boltzmann equation is directly connected with the electrical scattering mechanism Abrikosov88 ; Mahan81 . Although there is no consensus on the origin of the low-temperature electrical transport in the underdoped pseudogap phase to date, it is widely believed that the opening of the pseudogap induces a deviation from the low-temperature T-linear behaviour of the resistivity. In particular, it has been argued that if the electron scattering responsible for the low-temperature T-linear resistivity in the overdoped regime involves the electron scattering on the spin excitations in the underdoped regimePuchkov96 , then the opening of the spin pseudogap may naturally lead to a deviation from the low-temperature T-linear behaviour of the resistivityBucher93 ; Ito93 ; Nakano94 ; Ando01 . On the other hand, it has been shown clearly that the electron umklapp scattering is the origin of the low-temperature T-linear resistivity in the strange-metal phase Lee21 ; Honerkamp01 ; Hartnoll12 ; Ma23 . Moreover, weMa23 have shown very recently that the low-temperature T-linear resistivity in the overdoped strange-metal phase arises from the umklapp scattering between electrons by the exchange of the MF effective spin propagator, where the low-temperature electronic transport is mainly governed by the antinodal umklapp scattering, and then the low-temperature umklapp scattering rate is T-linear near the umklapp point, which naturally generates a low-temperature T-linear resistivity. In this paper, we employ the electron umklapp scattering to study the low-temperature electrical transport in the underdoped pseudogap phase by considering the spin pseudogap effects. To see the electron umklapp scattering process more clearly, we show a schematic picture of the electron umklapp scattering processLee21 ; Ma23 in Fig. 4, where an electron on a circular EFS (left) is scattered by its partner on the umklapp EFS (right). It should be noted that the intensity map of EFS in Fig. 4 is the same as the EFS map in Fig. 3, while the Fermi wave vector of the tips of the Fermi arcs ${\rm k}^{\rm TFA}_{\rm F}$ is the radius of the EFS circle (red). In particular, this circle EFS (red) connects all tips of the Fermi arcs, and then the most of the electronic density of states is concentred on this circular EFS.
@@
rather than the scattering between electrons via the emission and absorption of the spin excitationLee21 ; Ma23 .
In cuprate superconductors, a small density of charge carriers is sufficient to destroy the AF long-range order. However, the doped charge carriers and magnetic-exchange-coupled spins organize themselves in a cooperative way to enhance both the electrons mobility and the AF short-range order correlation, and then the spin excitations in the spin liquid state with the AF short-range order appear to survive from the underdoped regime to the overdoped regimeFujita12 . Moreover, it has been shown that the coupling strength of the electrons with the spin excitations gradually weakens with the increase of doping from a strong-coupling case in the underdoped regime to a weak-coupling side in the overdoped regime Feng12 ; Johnson01 ; Kordyuk10 , reflecting a reduction of the strength of the magnetic fluctuation with the increase of doping. In other words, (i) the effect of the magnetic fluctuation in the overdoped regime is less dramatic than in the underdoped regimeKastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; (ii) on the other hand, as shown in Fig. 2, the effect from antinodal spin pseudogap is absent in the overdoped regime, and then the main properties of the antinodal spin excitations can be well described by the MF spin propagator (9). These are reasons why the interpretation of the low-temperature T-linear resistivity in the overdoped strange-metal phase can be well made in terms of the umklapp scattering between electrons by the exchange of the MF effective spin propagatorMa23 .
However, in the strong-coupling side (then in the underdoped regime), (i) the effect of the magnetic fluctuation is much dramatic Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; and (ii) as shown in Fig. 2, the effect from antinodal spin pseudogap is particularly notable, where the antinodal spin pseudogap lowers the density of the spin excitations in response to the intense electron umklapp scattering, which reduces the strength of the electron umklapp scattering from the spin excitation states into the antinodal region. In this case, the umklapp scattering between electrons should be mediated by the exchange of the full effective spin propagator for a proper description of the low-temperature electrical transport in the underdoped pseudogap phase.
Now our aim is to obtain the full effective spin propagator. The full spin bubble $\bar{\Pi}({\bf p},{\bf q},\omega)$ in Eq. (32) is derived directly in terms of the full spin propagator (17) as,
@@
$\times$ $[n_{\rm B}(\bar{\omega}_{\alpha{\bf q}+{\bf p}})-n_{\rm B}(\bar{\omega}_{\alpha^{\prime}{\bf q}})].$
The full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (32) now can be derived from the above spin bubble (33) as,
@@
The electron umklapp scattering in Eq. (31 shows that the electron-electron collision $I_{\rm e-e}$ is both functions of momentum and energy. Howerver, the amplitude of the wave vector dependence is unimportant at low temperatures, since everything happens at EFSAbrikosov88 ; Mahan81 . In this case, an any given patch on the circular EFS shown in Fig. 4 is represented via the Fermi angle $\theta$ with the Fermi angle range $\theta\in[0,2\pi]$, and then the momentum integration along the perpendicular momentum can be replaced by the integrationLee21 ; Ma23 ; Prange64 over $\bar{\varepsilon}_{\bf k}$. For the umklapp scattering between electrons by the exchange of the full effective spin propagator in Eq. (31), an electron on the circular EFS parametrized by the Fermi angle $\theta$ is scattered to a point parametrized by the Fermi angle $\theta^{\prime}$ on the umklapp EFS via the spin excitation carrying momentum ${\rm p}(\theta,\theta^{\prime})$ as shown in Fig. 4. Following the above treatment and the calculation process in the recent worksLee21 ; Ma23 for the low-temperature electrical transport in the overdoped strange-metal phase, the electron-electron collision $I_{\rm e-e}$ in Eq. (31) in the underdoped pseudogap phase can be obtained straightforwardlyMa23 , and then the electrical Boltzmann transport equation (30) can be derived as,
@@
and then the electrical resistivity can be expressed directly in terms of the above dc conductivity as,

### v2

@@ II.2 Pseudogap in charge and spin channels
Based on the full spin propagator (8), the dynamical spin response of cuprate superconductors in the normal-state has been investigatedKuang15 ; Yuan01 ; Feng98 , where the obtained results show the existence of damped but well-defined dispersive spin excitations in the whole doping phase diagram. In particular, the spectral weight of the spin excitation spectrum at around the antinodal region is strongly suppressed by the antinodal spin pseudogapKuang15 . Moreover, the low-energy spin fluctuation is dominated by the process from the mobile charge carriers, while the high-energy spin excitation on the other hand retains roughly constant energy as a function of doping, with spectral weight and dispersion relation comparable to those in the corresponding SC-state. All these results are qualitatively consistent with the experimental observations Fujita12 ; Birgeneau89 ; Fong95 ; Yamada98 ; Arai99 ; Bourges00 ; He01 ; Tranquada04 ; Bourges05 .
@@ II.3 Electronic structure in the underdoped pseudogap state
For the understanding of the nature of the electronic structure in cuprate superconductors, it is needed to derive the electron propagator, which is characterized by the full charge-spin recombinationFeng15a . Following the previous discussionsFeng15a , the full electron propagator in the normal-state can be derived as [see Appendix A],
@@
where $\varepsilon_{\bf k}=-4t\gamma_{\bf k}+4t^{\prime}\gamma_{\bf k}^{\prime}+\mu$ is the electron energy dispersion in the tight-binding approximation, while the electron normal self-energy $\Sigma_{\rm ph}({\bf k},\omega)$ is given in Eq. (51) in Appendix A.
The electron normal self-energy in Eq. (11) originated from the charge-carrier normal self-energy in Eq. (3) is due to the charge-spin recombination Feng15a , indicating that the normal-state pseudogap state originated from the charge-carrier pseudogap state is due to the charge-spin recombination. To see the nature of the normal-state pseudogap more clearly, the above electron normal self-energy in Eq. (11) can be reexpressed as,
$\Sigma_{\rm ph}({\bf k},\omega)$$\approx$ ${[\bar{\Delta}_{\rm PG}({\bf k})]^{2}\over\omega-\varepsilon_{0{\bf k}}},$ (12)
where $\varepsilon_{0{\bf k}}=L_{2{\bf k}}/L_{1{\bf k}}$ is the energy spectrum of $\Sigma_{\rm ph}({\bf k},\omega)$, and $\bar{\Delta}^{2}_{\rm PG}({\bf k})=L^{2}_{2{\bf k}}/L_{1{\bf k}}$ is the normal-state pseudogap, with the functions $L_{1}({\bf k})=-\Sigma_{\rm pho}({\bf k},\omega=0)$ and $L_{2}({\bf k})=-\Sigma_{\rm ph}({\bf k},\omega=0)$, while the antisymmetric part $\Sigma_{\rm pho}({\bf k},\omega)$ of the electron normal self-energy can be obtained directly from the electron normal self-energy $\Sigma_{\rm ph}({\bf k},\omega)$ in Eq. (51). This normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ is also identified as being a region of the electron normal self-energy in which the normal-state pseudogap anisotropically suppresses the electronic density of states on EFS. Since the normal-state pseudogap state is induced by the charge-carrier pseudogap state, the normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ [the normal-state pseudogap crossover temperature $T^{*}$] as a function of doping presents a similar behavior of the charge-carrier pseudogap $\bar{\Delta}^{({\rm h})}_{\rm pg}({\bf k})$ [the charge-carrier pseudogap crossover temperature $T^{*}_{\rm h}$]Feng15 ; Feng12 . This normal-state pseudogap crossover temperature $T^{*}$ is actually a crossover line below which a novel electronic state emerges, as exemplified by the presence of the Fermi arcs, the competing electronic orders, etc., and then the unconventional features of this novel electronic state can be well interpreted in terms of the formation of the normal-state pseudogap Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 .
@@
$I({\bf k},\omega)\propto n_{\rm F}(\omega)A({\bf k},\omega),$ (13)
@@
where ${\rm Re}\Sigma_{\rm ph}({\bf k},\omega)$ and ${\rm Im}\Sigma_{\rm ph}({\bf k},\omega)$ are the real and imaginary parts of $\Sigma_{\rm ph}({\bf k},\omega)$, respectively. The electrons are renormalzed due to the electron scattering mediated by the spin excitation and then they acquire a finite lifetime.
The EFS topology is known to be crucial in the understanding of the unconventional superconductivity in cuprate superconductors Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 and the related anomalous normal-state properties Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 . The ARPES measured EFS can be described theoretically by the intensity map of the electron excitation spectrum (13) at zero energy $\omega=0$, where the locations of the EFS continuous contour in momentum space is determined directly by the poles of the electron spectral function (14), $\varepsilon_{\bf k}+{\rm Re}\Sigma_{\rm ph}({\bf k},0)=\bar{\varepsilon}_{\bf k}=0$, with the renormalized electron energy dispersion $\bar{\varepsilon}_{\bf k}=Z_{\rm F}\varepsilon_{\bf k}$ and the single-particle coherent weight $Z^{-1}_{\rm F}=1-{\rm Re}\Sigma_{\rm pho}({\bf k},0)\mid_{{\bf k}=[\pi,0]}$. However, the strong redistribution of the spectral weight on EFS is mainly governed by the momentum dependence of the normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ [then the single-electron scattering rate $\Gamma_{\bf k}(\omega)={\rm Im}\Sigma_{\rm ph}({\bf k},\omega)=\pi[\bar{\Delta}_{\rm PG}({\bf k})]^{2}\delta(\omega+\varepsilon_{0{\bf k}})]$.
@@
For a convenience in the following discussions of the resistivity in the underdoped pseudogap phase, the EFS mapFeng16 ; Liu21 ; Cao21 ; Zeng22 at doping $\delta=0.09$ with temperature $T=0.002J$ is replotted in Fig. 3, where the BZ center has been shifted by [$\pi,\pi$], and AN, TFA, and ND indicate the antinode, tip of the Fermi arc, and node, respectively. It thus shows clearly that the antinodal region of EFS becomes partially gapped, leading to that EFS consists, not of a closed contour, but only of four disconnected Fermi arcs centered around the nodes, in qualitative agreement with the experimental results Loeser96 ; Norman98 ; Chatterjee06 ; Shi08 ; Sassa11 ; Kaminski15 ; Loret17 ; Loret18 ; He14 ; Comin14 ; Comin16 . However, the most of the spectral weight on the Fermi arcs locates at around the tips of the Fermi arcs, indicating that the electrons at around the tips of the Fermi arcs have a largest density of states Chatterjee06 ; Shi08 ; Sassa11 ; Kaminski15 ; Loret17 ; Loret18 ; He14 ; Comin14 ; Comin16 , and then charge order is driven by this EFS instability, with a characteristic wave vector corresponding to the tips of the Fermi arcsFeng16 ; Comin14 ; Comin16 .
The elementary excitations are parameterized by the electron spectral function (14), which has given a consistent description of the renormalization of the electrons in the underdoped pseudogap phaseFeng16 ; Liu21 ; Cao21 ; Zeng22 , where the obtained results show that the single-electron scattering rate has a well-pronounced peak structure at around the antinodal and nodal regions, which leads to the remarkable peak-dip-hump structure in the energy distribution curve. Moreover, the dispersion kink is induced by the inflection point in the single-electron scattering rate, while the spectral weight at around the dispersion kink is reduced highly by the corresponding peak in the real part of the electron normal self-energy. All these results are in qualitative agreement with the corresponding experimental results Dessau91 ; Randeria95 ; Fedorov99 ; Campuzano99 ; DMou17 ; Bogdanov00 ; Kaminski01 ; Johnson01 ; Iwasawa08 ; Plumb13 . More importantly, the anisotropic suppression of the electronic density of states on EFS by the normal-state pseudogap can affect the electrical transport in two ways Timusk99 : (i) through the reduction of the number of electron current-carrying states as a normal-state pseudogap forms; and (ii) since the electron current carriers are scattered by the spin excitation, through the reduction in the density of spin excitations.
@@ II.4 Electrical transport due to umklapp scattering from a spin excitation
We now turn to discuss the low-temperature electrical transport in the underdoped pseudogap phase. Theoretically, the Boltzmann transport equation is the cornerstone for the discussions of the electrical transportAbrikosov88 ; Mahan81 , since the Boltzmann transport equation is valid in the case of either the existence of the well-defined quasiparticles or the treatment of the electron interaction mediated by different bosonic modes within the Eliashberg approach. This builds on the following pioneering works: (i) in the early days of the electrical transport research in the conventional superconductorsPrange64 , Prange and Kadanoff demonstrated that in an electron-phonon system, a set of transport equations can be derived in the Migdal’s approximation, where the electron interaction mediated by the phonon leads to the electron self-energy and vertex correction. In particular, this coupled set of transport equations for the electron and phonon distribution functions is correct even in the case of the absence of the well-defined quasiparticlesPrange64 . Nevertheless, one of the forms of the electrical transport equation,
@@
In the following discussions, we study the the low-temperature resistivity in the underdoped pseudogap phase based on the Boltzmann transport equation (15). For the calculation of this Boltzmann transport equation (15), the linear perturbation from the equilibrium in terms of the distribution function has been introduced Lee21 ; Ma23 ; Prange64 ,
@@
with the fermion distribution functions $n_{\rm F}(\omega)$, and the local shift of the chemical potential at a given patch of EFS $\tilde{\Phi}({\bf k})$, which satisfies the antisymmetric relation $\tilde{\Phi}(-{\bf k})=-\tilde{\Phi}({\bf k})$. We substitute the above result in Eq. (16) into Eq. (15), and then linearize the Boltzmann equation (15) as,
@@
The electron-electron collision term $I_{\rm e-e}$ in the Boltzmann equation is directly connected with the electrical scattering mechanism Abrikosov88 ; Mahan81 . Although there is no consensus on the origin of the electrical transport in the underdoped regime to date, it is widely believed that the opening of the spin pseudogap induces a deviation from the T-linear behaviour of the resistivityGurvitch87 ; Bucher93 ; Ito93 ; Nakano94 ; Ando01 . In particular, it has been shown clearly that the electron umklapp scattering is the origin of the low-temperature T-linear resistivity in the strange-metal phase Lee21 ; Honerkamp01 ; Hartnoll12 . Moreover, weMa23 have also shown very recently that the low-temperature T-linear resistivity in the overdoped strange-metal phase originates from the umklapp scattering between electrons by the exchange of the MF effective spin propagator. In the following discussions, we will show that when this electron umklapp scattering mediated by the spin excitation in the overdoped regimeMa23 flows to the underdoped regime, the opening of the momentum dependence of the spin pseudogap naturally leads to a low-temperature T-quadratic behaviour of the resistivity. To see the electron umklapp scattering process more clearly, we show a schematic picture of the electron umklapp scattering processLee21 ; Ma23 in Fig. 4, where an electron on a circular EFS (left) is scattered by its partner on the umklapp EFS (right). It should be noted that the intensity map of EFS in Fig. 4 is the same as the EFS map in Fig. 3, while the Fermi wave vector of the tips of the Fermi arcs ${\rm k}^{\rm TFA}_{\rm F}$ is the radius of the EFS circle (red). Moreover, this circle EFS (red) connects all tips of the Fermi arcs, and then the most of the electronic density of states is concentred on this circular EFS.
@@
rather than the scattering between electrons via the emission and absorption of the spin excitationLee21 ; Ma23 , where $\Lambda_{{\bf k}}=4t\gamma_{\bf k}-4t^{\prime}\gamma_{\bf k}^{\prime}$ is the bare vertex function.
In cuprate superconductors, a small density of charge carriers is sufficient to destroy the AF long-range order (AFLRO). However, the doped charge carriers and the coupled spins organize themselves in a cooperative way to enhance both the electrons mobility and the AF short-range order (AFSRO) correlation, and then the spin excitations in the spin liquid state with AFSRO appear to survive from the underdoped regime to the overdoped regimeFujita12 . Moreover, it has been shown that the coupling strength of the electrons with the spin excitations gradually weakens with the increase of doping from a strong-coupling case in the underdoped regime to a weak-coupling side in the overdoped regime Feng12 ; Johnson01 ; Kordyuk10 , reflecting a reduction of the strength of the magnetic fluctuation with the increase of doping. In other words, (i) the effect of the magnetic fluctuation in the overdoped regime is less dramatic than in the underdoped regimeKastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; (ii) on the other hand, as shown in Fig. 2, the effect from antinodal spin pseudogap is absent in the overdoped regime, and then the main properties of the antinodal spin excitations can be well described by the MF spin propagator. These are reasons why the interpretation of the low-temperature T-linear resistivity in the overdoped strange-metal phase can be well made in terms of the umklapp scattering between electrons by the exchange of the MF effective spin propagatorMa23 .
However, in the strong-coupling side (then in the underdoped regime), (i) the effect of the magnetic fluctuation is much dramatic Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; and (ii) as shown in Fig. 2, the effect from antinodal spin pseudogap is particularly notable, where the antinodal spin pseudogap lowers the density of the spin excitationsKuang15 in response to the intense electron umklapp scattering, which reduces the strength of the electron umklapp scattering from the spin excitation states into the antinodal region. In this case, the umklapp scattering between electrons should be mediated by the exchange of the full effective spin propagator for a proper description of the low-temperature resistivity in the underdoped pseudogap phase.
Now our aim is to obtain the full effective spin propagator. The full spin bubble $\bar{\Pi}({\bf p},{\bf q},\omega)$ in Eq. (19) is a convolution of two full spin propagators, and can be derived directly from the full spin propagator (8) as,
@@
$\times$ $[n_{\rm B}(\bar{\omega}_{\alpha{\bf q}+{\bf p}})-n_{\rm B}(\bar{\omega}_{\alpha^{\prime}{\bf q}})],$
where $n_{\rm B}(\omega)$ is the boson distribution function. From the above spin bubble (20), the full effective spin propagator (19) now can be derived as,
@@
The electron umklapp scattering in Eq. (18) shows that the electron-electron collision $I_{\rm e-e}$ is both functions of momentum and energy. Howerver, at low temperatures, everything happens near EFSAbrikosov88 ; Mahan81 . In this case, an any given patch on the circular EFS shown in Fig. 4 is represented via the Fermi angle $\theta$ with the Fermi angle range $\theta\in[0,2\pi]$, and then the momentum integration along the perpendicular momentum can be replaced by the integrationLee21 ; Ma23 ; Prange64 over $\bar{\varepsilon}_{\bf k}$. For the umklapp scattering between electrons by the exchange of the full effective spin propagator in Eq. (18), an electron on the circular EFS parametrized by the Fermi angle $\theta$ is scattered to a point parametrized by the Fermi angle $\theta^{\prime}$ on the umklapp EFS via the spin excitation carrying momentum ${\rm p}(\theta,\theta^{\prime})$ as shown in Fig. 4. Following the above treatment and the calculation process in the recent worksLee21 ; Ma23 for the low-temperature resistivity in the overdoped strange-metal phase, the electron-electron collision $I_{\rm e-e}$ in Eq. (18) in the underdoped pseudogap phase can be obtained straightforwardlyMa23 , and then the Boltzmann transport equation (17) can be derived as,
@@
At low temperatures, the electron elastic scattering occurs on EFS, while the electron inelastic scattering occurs near EFSKohsaka08 ; Abrikosov88 ; Mahan81 . In this case, the above energy integration in the kernel function (28) actually includes the inelastic scattering process of momentum and energy exchange between electronsLee21 ; Ma23 ; Prange64 . This follows from a fact that the umklapp scattering rate $\gamma(\theta)$ in Eq. (27) is directly associated with the kernel function $F(\theta,\theta^{\prime})$ in Eq. (28), while this kernel function is obtained by the integration of the spin excitation energy $\omega$, leading to the occurrence of the inelastic scattering process near EFS in the present umklapp scattering. In particular, in the case of the absence of the spin pseudogap, this electron umklapp scattering induces a low-temperature T-linear resistivity in the overdoped regimeMa23 , in agreement with the experimental observations Legros19 ; Ayres21 ; Grisso21 , where the experimental analyses seem to indicate that the T-linear resistivity is due to the inelastic scattering.
@@
and then the resistivity can be expressed directly in terms of the above dc conductivity as,